OpenStudy (anonymous):

What is the solution of the following system? -2x-y=1 -4x-2y=-1

OpenStudy (anonymous):

A. Infinitely many solutions B. No solution c. (3,8) D. (-3, -8)

mathslover (mathslover):

It can be solved by 2 methods : i) Substitution ii) Elimination . Do you know both of these methods? If yes, then which one you want to solve with.

mathslover (mathslover):

Ok, good question. No need to solve this. See, do you know the formula for checking whether there will be no solution, infinitely many solutions or some solution...

OpenStudy (anonymous):

Uhm, No I do not..

mathslover (mathslover):

Ok, suppose I have two equations: \(\color{blue}{\large a_1 x + b_1 y = c_1}\\ \large{\color{red}{a_2 x + b_2 y = c_2 } }\) Then: \(\large \color{blue}{i) \space \textbf{If :} \cfrac{a_1}{a_2} \ne \cfrac{b_1}{b_2} \ne \cfrac{c_1}{c_2} \implies \textbf{The equations will have some solutions}} \\ \color{red}{ii) \space \textbf{If : } \cfrac{a_1}{a_2} = \cfrac{b_1}{b_2} \ne \cfrac{c_1}{c_2} \implies \textbf{The equations will have no solution }} \\ \color{orange}{iii) \space \textbf{If :} \cfrac{a_1}{a_2} = \cfrac{b_1}{b_2} = \cfrac{c_1}{c_2} \implies \textbf{The equations will have infinitely many solutions}} \)

mathslover (mathslover):

Please notice the formula again, I had edited the mistake there. Sorry for inconvenience Now, in : -2x - y = 1 , can you tell me what is \(a_1\) , \(b_1\) , \(c_1\) ... ? by comparing it with \(a_1x+b_1 y = c_1\)

mathslover (mathslover):

@SamiiBelle ?

OpenStudy (anonymous):

Thank you for this. I worked the problem out with my dad and I got No solutions.

mathslover (mathslover):

Good work, yes it is No Solutions. :) Give a "thumbs up" to your dad too :)

OpenStudy (anonymous):

Thanks for the metal! @mathslover